Mission Support COMMS closed 23Apr2024

Mission support is signing off.
Please see below received reports status:

Sol Summary Received
Journalist Report Received
GreenHab Report Received
Operations Report Received
EVA Report Received (2)
Photos (6-8 pics) Received
EVA Request Approved

Ben Stanley – Site ManagerMars Desert Research Station



Journalist Report – April 20th

Crew 297 – Janus 1 Journalist – David Laude

As the Martian sun slowly cast its crimson glow over the horizon, the crew of the Janus 1 awoke to the start of another extraordinary day amidst the barren yet captivating landscape of the Red Planet. With near ideal weather conditions, the EVA team departed the Hab at 9:08 AM and mounted their all electric rovers, Opportunity and Curiosity. The EVA participants were Pawel Sawicki (Commander), Sarah Lamm (Geologist), Matthew Lynch (Engineer), Matthew Storch (Executive Officer and EVA Team Leader)

Team leader Matt S. remarked in detail: "We first did another seek for the NPS Project, by traveling to the coordinates specified by Mission Support, and searching a 300 m radius for the NPS. We then evaluated the robot, excavated a site for placement of the NPS, and explored the surrounding region for geological purposes."

"Crew obtained the NPS search coordinates and proceeded to the search area, which was near the Candor Chasma. We stopped just off Cow Dung road and proceeded on foot along the Gateway to Candor, to conduct the search. The crew split into 2 teams that straddled a ravine. Matt Lynch quickly spotted the NPS and we retrieved it and returned to the rovers. We then proceeded north along Cow Dung Road to Galileo Road and headed east on Galileo until we reached the turnaround limit on battery power. The turnaround point was not deemed suitable for the other mission objectives, so we started to return along Galileo road, stopping twice to evaluate possible sites to cover the remaining mission objectives. Once a suitable location was found, Pawel and Matt S set up the robot, while Matt L scouted for suitable nearby excavation sites and Sarah conducted geological studies of the area. The robot had problems connecting to its network and there was not much we could do in the field to address the problem, so we had to abandon that objective for today. All four EVA crew members then joined the excavation effort, using a variety of tools and working in shifts. In the end we were able to excavate deep enough to properly bury the NPS, but finished just in time to return from the EVA with a suitable time margin. Excavation was much more difficult than in the first location, with the regolith removal rate being about one half of what it was in the first excavation, despite having 4 crew members to work this excavation vs. 2 for the previous one."

Meanwhile, at the Hab, Dave was preparing his first scone with blueberries mix for the oven. Despite the use of freeze dried foods for most of the ingredients, the scones came out of the oven in fine shape and satisfying taste. The fine shape was like a flower pedal.

Today’s featured crew member is Pawel. In his biography he notes that he "is currently a New Shepard Crew Capsule Test Engineer at Blue Origin, where he is responsible for the successful and safe execution and on-time completion of several major launch vehicle tests and pre-flight checks. At Blue Origin, he is also a volunteer Emergency Response Team member. Pawel earned a Ph.D. from the University of Colorado at Boulder, where his doctoral research involved computationally investigating amelioration techniques for plasma-induced radio wave blackout, which has historically plagued hypersonic vehicles. Pawel had also obtained an M.S. in Biomedical Engineering from the University of Colorado at Boulder, an M.S. in Aerospace Engineering from the University of Michigan, and M.S. and B.S. degrees in Mechanical Engineering from New York University. Pawel’s career has also included stints of varying capacities at NASA Ames Research Center, NASA Langley Research Center, NASA Marshall Space Flight Center, and Lockheed Martin Advanced Technology Center."

Sol Summary – April 20th

Crew 297 Sol Summary Report 19-04-2024

Sol: 06

Summary Title: Live Another Sol

Author’s name: Pawel Sawicki

Mission Status: Nominal

Sol Activity Summary:

Our Sol 6 went quite a bit smoother than the infamous Sol 6 of Ares III and Mark Watney.

The crew started the sol feeling rejuvenated after some great sleep, induced by indulging on the delicious late-night apple pie made by Matt Lynch. Shortly after waking up in the morning, the crew started getting ready for their 4-hour EVA. They were provided coordinates of the desired landing spot for the NPS, though strong winds last night increased the NPS search radius to 400 meters. Fortunately for the crew, it took only several minutes to find the NPS and carry it back to the rovers. This left more than three hours for the crew to pursue objectives for their respective research projects. Sarah utilized the additional time to log gamma-ray spectral data of samples within the beautiful landscape. The Case robot unfortunately was quite shy today, and did not want to communicate with the rest of the crew, consequently shelving his objective for another day. "I guess you could call it a failure," but I prefer the term "learning experience" (Watney, in The Martian, Weir). With an additional crewmate available for excavating, a hole adequate for the NPS was dug in 67 minutes, even through a challenging surface.

The crew came back from their EVA to freshly-made scones, prepared by Dave. The rest of the afternoon was spent performing nominal duties, conducting and preparing research, and writing reports. Dinner was made by Chef Marquez, a refined shoyu ramen recipe exquisitely composed of a dashi-base crafted with kombu seaweed, shiitake mushroom, and tare. The ramen was then brought together with an egg-based ramen, kanpyo, flavored tomato, onion, and GreenHab-harvested red cabbage and green onion. The umami-packed, harmonious mixture came together to bring us all closer together in the later hours of the evening.

Look Ahead Plan:

Tomorrow will be a rest day of sorts for the crew with a late wake-up and less work planned than usual. Nevertheless, there are two low-activity EVAs planned tomorrow. The first EVA will involve Team Matt picking a suitable location for the NPS and hiding it for a second team to recover. The second EVA will be more focused around the station, specifically at the RAM. Given Case’s difficulties today, the robot will be tested right outside of the Hab. Dave will manually drive the robot from a ground station in the RAM while Pawel monitors closeby.

Anomalies in work:

Suit 11 (fan power significantly decreased mid-EVA): In work
GreenHab automated fan: Work-around has been implemented to power on the fan manually. In case of power outage or inoperable fan, the GHO will open the door for cooling.
ScienceDome: We have noticed a foul chemical smell the last few days which has made the ScienceDome unsuitable for any work.
EVA headset #2 not receiving comms when connected with two way radio: R&R with spare headset.
Morning (data from 0900): calm winds (2 mph), chilly (57F), and sunny with clear skies
Noon: calm winds (2 mph), warm (68F), and few clouds in the sky
Afternoon (1800): moderate winds (19 mph), warm (74F), and sunny with clear skies
Crew Physical Status: EVA Crew members are physically tired from the EVA today, but recovering well. Overall crew health is nominal.

EVA: Today’s EVA was nominal and an overall success. The crew of 4 (Sawicki, Storch, Lynch, Lamm) was able to locate the hidden NPS quickly today, within 10 minutes of parking the rovers and taking off on foot. The crew continued with their secondary mission objectives. Gamma-ray spectral data was collected from samples in the area and the crew was able to dig (and fill) a hole suitable for the NPS. Unfortunately, Case had network issues and was not utilized today.

Reports to be filed:

Sol Summary
Journalist Report
GreenHab Report
Operations Report
EVA report (1 EVA)
EVA request (2 EVAs)
Mid-Mission Report
Support Requested: None requested.

Journalist Report – April 22nd

Crew 297 – Janus 1 Journalist – David Laude

"Mars is there, waiting to be reached." – Buzz Aldrin

EVA 11 team, Pawel and Sarah, did some reaching to Mars in sim today, driving as far as rover Oppy would take them and still make it back to base on one charge. They arrived at the grey unit and took gamma-ray measurements of it, looking for uranium content and comparing to previously measured units. Later they walked to the overlook and then to south of the Sea of Shells. Finally, they hid the NPS for tomorrow’s morning EVA team to search for. Those remaining at the MDRS campus worked on a variety of odds and ends throughout the day.

Today’s featured crew member is Sean Marquez, our GreenHab Officer. In his biography he states: "Sean has a B.S. degree in Mechanical Engineering, specializing in design of mechanical systems, from the University of California, Irvine. He worked as an associate mechanical design engineer for Max Q Systems – formerly an original equipment manufacturer (OEM) for the aerospace industry. In his spare time, he contributes to FPrime, an open-source flight software and embedded systems framework used on the NASA/JPL Mars Ingenuity helicopter and university CubeSATs. FPrime is currently undergoing implementation at the Mars Desert Research Station to automate monitoring of plants in the GreenHab. Sean also works with a working group with the Open Source Hardware Association (OSHWA) and the Mach 30 Foundation to develop open standards for the medical/aerospace industry, as well as methodologies for developing open-source hardware (OSHW) like open-source software (OSS). Sean is currently studying permaculture design to develop a means to becoming multiplanetary without the need for interplanetary supply chains, using permaculture as its guiding principles."

Operations Report – April 22nd

Crew 297 Operations Report 22-04-2024

SOL: 8

Name of person filing report: Matt Lynch (Engineer)

Non-nominal systems: Suit 2 (stuck valve), Suit 9 (broken visor), Suit 11 (fan issue)

Notes on non-nominal systems: The issues with Suit 2 and 9 were already noted before our mission. Suit 11’s battery was replaced yesterday, but on today’s EVA Suit 11 was believed to exhibit reduced circulation (albeit better than before the battery was replaced). After inspection the cause remains unknown. 3 missing zip ties for the tarps in the tunnels were replaced.


Opportunity rover used:

Hours: 183.4

Beginning charge: 100%

Ending charge: 46%

Currently charging: No

General notes on rovers: NA

Summary of Hab operations:

WATER USE: 25.4 gallons

Water (static tank): 233.4 gallons

Static tank pipe heater (on or off): off

Static tank heater (On or off): off

Toilet tank emptied: no

Summary of internet: Looking up recipes, downloading entertainment, catching up on Earthly news

Summary of suits and radios:

Radios 1,5, and 6 were used.

EVA Suits 4 and 7 were used.

Summary of GreenHab operations:

WATER USE: 4 gallons

Heater: Yes

Supplemental light: N/A

Harvest: 2g of cherry tomatoes

Summary of ScienceDome operations: Setup for Pawel’s EVA tomorrow took place in the ScienceDome now that the noxious smell is gone.

Summary of RAM operations: (Tools used, work done): 3 missing zip-ties for the tarps on the tunnels were replaced using the supplies from the RAM.

Summary of any observatory issues: NA

GreenHab Report – April 22nd

Crew 297 GreenHab Report 22-04-2024

GreenHab Officer: Sean Marquez

Environmental control:

Heater on (set to 55F)
Fan on (set to turn on at 94F; turn off at 82F)
Average temperatures: 82F

Hours of supplemental light: N/A

Daily water usage for crops: 4 gal

Daily water usage for research and/or other purposes: 0 gal

Water in Blue Tank: 175 gal

Time(s) of watering for crops: 12:50

Changes to crops: N/A

Narrative: Fan system was operating nominal. Though, due to the afternoon heat wave, the GreenHab atmospheric temperature reached close to 110F around 14:30. The GreenHab door was opened again to allow for supplemental airflow from the tunnel system, which helped keep the atmospheric temperatures below 100F. More brown/yellow leaves/stems from the cherry tomato plants were composted.


2g of cherry tomato
Support/supplies needed: N/A

EVA Report – April 22nd

Crew 297 EVA 11 Report 21-4-2024    

EVA # 11

Author: Sarah Lamm

Purpose of EVA: To observe the grey geological unit seen from orbit.

Start Time: 7:37 AM

End Time: 10:11 AM

Narrative: We arrived at the grey unit and took gamma-ray measurements of it, specifically to see if the uranium content would be higher than the previous units. From there we traveled to the overlook, and just south of the Sea of Shells. Once we came back to the intersection of Brahe Hey and Cow Dung Row, we hid the NPS device for Matt’s L project.

Destination: Area between Sea of Shells and Overlook

Coordinates: 12S 516033 4254592

Participants: Pawel Sawicki (Commander), Sarah Lamm (Geologist)

Road(s) and routes per MDRS Map: Cow Dung Road, Brahe Hwy

Mode of travel: Opportunity

Sol Summary – April 21st

Crew 297 Sol Summary Report 21-04-2024

Sol: 07

Summary Title: I’m Sorry Dave, I’m Afraid I Can’t Do That

Author’s name: Pawel Sawicki

Mission Status: Nominal

Sol Activity Summary:

The crew woke up today well rested after a night of watching the Martian and getting to sleep in, a celebration for passing the mid-mission mark. Today’s docket of EVAs consisted of i) the Matts adventuring far northwest searching for an NPS hiding spot around the Sea of Shells and ii) Case adventuring right outside of the RAM. Both EVA crews were satisfied with what they saw today, whether it was the spectacular views of the Overlook or a fully functioning User Interface. The whole crew was later treated to a bon gou Jambalaya by Chef Lamm, a dish reminiscent of New Orleans street jazz – soulful, vibrant, and complex characteristics heard within each bite.

Look Ahead Plan:

Tomorrow there will only be one EVA, commencing shortly after the Martian sun wakes up. During this EVA, Sarah and Pawel intend on exploring interesting geological features to the northwest of the Hab within the vicinity of the Sea of Shells and Overlook. The rest of the crew will rest from EVAs and continue their nominal duties including preparing and conducting IVA research.

Anomalies in work:

Suit 11 (fan power significantly decreased mid-EVA): Battery has been replaced.
GreenHab (automated fan not working): working now that generator power is back online
ScienceDome (foul smell): Source of odor (batteries) has been determined and removed. Crew will continue to monitor for another day before using the ScienceDome fully.
Morning (data from 0900): calm winds (2 mph), chilly (57F), and sunny with clear skies
Noon: calm winds (2 mph), warm (68F), and sunny with clear skies
Afternoon (1800): moderate winds (14 mph), warm (77F), and sunny with clear skies
Crew Physical Status: Crew health is nominal.

EVA: There were two EVAs today with two crewmates each – a first for us. The first EVA crew set out to hide the NPS, but due to a lack of good hiding spots in the area this goal was delayed to a later EVA. The crew remarked on the spectacular landscapes that part of Mars has to offer. The other EVA was more local to the Hab (by the RAM), where Case went out for a walk to test its features. Operation of the robot was as smooth as expected and more challenging EVAs are to come for Case.

Reports to be filed:

Sol Summary
Journalist Report
GreenHab Report
Operations Report
EVA report (2 EVAs)
EVA request (1 EVA)

Support Requested: None requested.

Copyright © The Mars Society. All rights reserved. | Main Site